OpenStudy (anonymous):

Given the function g(x) = 8x − 2, find g(−2).

OpenStudy (turingtest):

g(-2)=8(-2)-2

OpenStudy (anonymous):

plug in 2 ...

OpenStudy (anonymous):

g(-2) = 8(-2) - 2 g(-2) = -16 - 2 g(-2) = -18

OpenStudy (anonymous):

8*(-2)-2=-18